OpenStudy (surryyy):

The perimeter of a rectangle is one more than the area of the rectangle. Find the area of the rectangle, if its width is 3 cm. Answer 5 cm2 16 cm2 225 cm2 15 cm2

OpenStudy (anonymous):

A = width x length P= 2(width +length) 1 + Area = Perimeter Width = 3 I believe the answer is 15cm2

OpenStudy (shinalcantara):

Formulas for the area and perimeter of a rectangle A = LW ----equation 1 P = 2L + 2W ----equation 2 -------------------- Given that the perimeter is 1 more than the area, you'll have the expression as: A + 1 = P ----equation 3 and that W=3cm -------------------- Expressing equation 3 in terms of L and W by substituting equations 1 and 2 you'll have: A + 1 = P (LW) + 1 = (2L + 2W) LW + 1 = 2L + 2W Since W=3cm, then we'll substitute to find the value of L L(3) + 1 = 2L + 2(3) 3L + 1 = 2L + 6 Transpose 2L to the left side and 1 to the right side 3L - 2L = 6 - 1 Simplify L = 5 -------------------- To find the area, go back to equation 1 A = LW W = 3 and L = 5 A= 5cm(3cm) A = 15cm^2